### v2

@@ Abstract
The reconstruction of interaction networks between random events is a critical problem arising from statistical physics and politics, sociology, biology, psychology, and beyond. The Ising model lays the foundation for this reconstruction process, but finding the underlying Ising model from the least amount of observed samples in a computationally efficient manner has been historically challenging for half a century. Using sparsity learning, we present an approach named SLIDE whose sample complexity is globally optimal. Furthermore, a tuning-free algorithm is developed to give a statistically consistent solution of SLIDE in polynomial time with high probability. On extensive benchmarked cases, the SLIDE approach demonstrates dominant performance in reconstructing underlying Ising models, confirming its superior statistical properties. The application on the U.S. senators voting in the last six congresses reveals that both the Republicans and Democrats noticeably assemble in each congress; interestingly, the assembling of Democrats is particularly pronounced in the latest congress.
Keywords: Ising model, best-subset selection, pseudo-likelihood, polynomial time complexity, consistent algorithm
@@ 1 Introduction
In contemporary applications, the Ising model serves as a prevalent tool for characterizing the pairwise interactions between binary variables in a network system. In this network, each node represents a binary variable, the presence of an edge between any two nodes indicates that the corresponding variables have an interaction. Reconstructing the corresponding graph of the Ising model is critical to the studies of interaction in a broad spectrum of complex systems in biology, neuroscience, ecology, economics, political science, image analysis (Cocco et al., 2009; Weistuch et al., 2021; Noble et al., 2018; Battiston et al., 2016; Banerjee et al., 2008; Saremi and Sejnowski, 2013). While many methods have been proposed during the past half century for reconstructing Ising models, finding the most sample-efficient estimate in a computationally efficient manner such as in a polynomial time remains an open problem. Given the profound importance of the Ising model, especially in the era of social media, there is an urgent need to solve this open problem.
@@ 1.1 Literature Review
Reconstructing Ising models by maximizing the likelihood function is considered to be the gold standard, although it is computationally intractable since the evaluation of the partition function involves the sum of $2^{p}$ terms (Nguyen et al., 2017). Many techniques have been developed to approximate the likelihood function. For instance, it is suggested that Monte Carlo methods can be used to approximate the partition function (Geyer, 1994; Miasojedow and Rejchel, 2018), and then the approximating likelihood function is maximized. Note that although this approach is asymptotically exact, it takes an exponential number of runs to achieve a predefined accuracy. Alternative methods that can avoid evaluating the partition function include the naive mean-field approximation and its advanced variants (Chow and Liu, 1968; Kappen and de Borja Rodríguez Ortiz, 1997; Nguyen and Berg, 2012). These variational methods assume that the network is composed of some basic clusters, and when this assumption is met, they give the exact value of the likelihood function. As one increases the size of basic clusters, the methods apply to a broader class of networks, while the computational complexity also shapely grows. However, for networks beyond the applicable class, the effectiveness of such approximations remains un-investigated.
@@
Instead of approximations, surrogates of the likelihood function also have been proposed to sidestep the computational cost. Ravikumar et al. (2010) proposed to node-wisely maximize the conditional likelihood. The product of conditionals, commonly referred to as pseudo-likelihood or composite likelihood, is often coupled with various regularizations and widely applied in the literature (Besag, 1975; Höfling and Tibshirani, 2009; Jalali et al., 2011; Xue et al., 2012). In Vuffray et al. (2016), one proposed to use the interaction screening objective, which is closely related to the exponential loss of the logistic model. These methods are capable for general Ising models and can alleviate the prohibitive computation for their reconstruction. Disadvantages of this strategy include estimation biasedness, as well as the subtle steps in tuning hyperparameters and selecting post-inference thresholds for theoretic guarantees (Lokhov et al., 2018). More importantly, we prefer methods that use the fewest samples to perfectly recover the underlying Ising model. However, every aforementioned method requires sample sizes much larger than the necessary sample complexity derived in Santhanam and Wainwright (2012). The best result is the optimality along only a subset of the parameters (Lokhov et al., 2018).
@@
Graphical structure recovery resides in the realm of feature selection, for which the predominance of sparse learning approach has been demonstrated in a variety of statistical models (Raskutti et al., 2011; Sara van de Geer and Peter Bühlmann, 2013). The sparse learning approach directly addresses the biasedness of the aforementioned regularization-based methods. Moreover, the approach may lead to stronger statistical properties, as they usually result in a smaller parameter space to search over. Finally, the development of statistical tools allows us to consider directly solving the sparse learning problem. For some well-studied models like (generalized) linear models, in recent years, statisticians have proposed several feasible approximate algorithms that solve the sparse regression under mild conditions (Huang et al., 2018; Zhu et al., 2020; Zhang et al., 2023; Zhu et al., 2022). However, for Ising models, the potential benefits of the sparse learning approach have not yet been rigorously authenticated, and how to design a feasible algorithm that captures the essence of sparse learning for reconstructing the Ising model remains another meaningful problem.
@@ 2.1 Ising Model and its Reconstruction
Definition 1 (Family of distributions, Santhanam and Wainwright (2012))
@@
$P_{J}(z)=\frac{1}{\Phi({J})}\exp\Big{(}\sum_{i<j}{J}_{ij}{}_{i}{}_{j}\Big{)}$
since ${J}_{ij}=0$ implies $(i,j)\notin E$. Note that, we occasionally omit the subscript ${J}$ in $P_{{J}}$ for simplicity when there is no ambiguity; besides, we define the support set of matrix ${J}$ as ${\mathrm{supp}}({J})$.
Suppose ${J}^{*}$ is underlying but unknown that characterizes variables’ interactions, Ising model reconstruction aims to recover ${J}^{*}$ via $n$ observed samples $z^{(1)},\ldots,z^{(n)}\in\mathbb{R}^{p}$ that are independently sampled from the Ising model $P_{{J}^{*}}$. Specifically, in this paper, we focus on this problem when the underlying graph only has a relatively small number of edges, i.e., the vast majority of the ${J}_{ij}$’s are exactly zero. This sparsity assumption is reasonable as it gives better interpretability of the model. In the meanwhile, it appears to be a necessary assumption when the dimension $p$ is comparable to or larger than the sample size $n$ (so-called the high-dimensional regime).
@@ 2.2 Globally Optimal Sample Complexity Regime
Background. For the reconstruction of Ising models, the sample complexity of an estimator refers to the minimum sample size such that, with a high probability, this estimator can recover any underlying Ising model belongs to the family of distributions introduced in Definition 1. For Ising models satisfying Definition 1, Santhanam and Wainwright (2012) derived the information-theoretic lower bound of the sample size with respect to four structural parameters: maximum neighborhood weight ${\gamma}$, maximum node degree $d$, number of nodes $p$, and minimum signal $\lambda$. These parameters are of primary interest when studying the sample complexity. Specifically, Santhanam and Wainwright (2012) showed that information-theoretic lower bound is of order
@@
Moreover, to show the tightness of the above lower bound, Santhanam and Wainwright (2012) provided an attainable sample complexity, by analyzing the sample complexity of a projection-type graphical structure estimator. To be specific, this estimator returns the graphical structure of the model whose second-order moments best match their empirical counterparts. Such an estimator is computationally intractable, but it offers a sample complexity that is attainable (at least theoretically) by an estimator:
@@
The optimal sample size $n_{opt}$ lies within $n_{l}$ and $n_{u}$ yet its specific form remains unknown. From $n_{l}$ and $n_{u}$, we know that the optimal sample complexity $n_{\text{opt}}$ has the following dependencies on $({\lambda},{\gamma},d,p)$: $n_{\text{opt}}\propto\log p$, $n_{\text{opt}}\propto{\lambda}^{-2}$, $n_{\text{opt}}\propto d^{r}$ and $n_{\text{opt}}\propto e^{c{\gamma}}$ where $r$ and $c$ are two constants satisfying $r\in[1,2]$ and $c\in[1,4]$.
@@
for any $c\in[1,4]$. An estimator whose sample complexity satisfies (3) is called a globally optimal estimator. The regime (3) integrates insights from both the information-theoretic lower bounds (1) and the attainable sample complexities (2) derived by Santhanam and Wainwright (2012). By defining the regime as the lower and upper limits for each critical structural parameter, this regime comprehensively covers the entire range of these parameters. The rationality of the globally optimal regime (3) is twofold.
  • Tightness. To the best of our knowledge, no existing algorithm has a lower sample complexity than our upper boundary, showing this regime as the tightest one that contains the optimal rate in the literature. Additionally, this regime incorporates the conjecture by Santhanam and Wainwright (2012) that the information-theoretic lower bound for $d$ is tight, further reinforcing the tightness of the defined regime.
@@
    overlooked the dependency on the maximum degree $d$. Their regime implicitly assumes $d=o(e^{\gamma})$, where the effect of $d$ in both of (1) and (2) can be ignored and they merely focus on the effect of ${\gamma}$. However, this assumption may not hold for general Ising models and a counterexample is constructed in Section 5.1.1. By removing the assumption, the globally optimal regime serves as a more comprehensive extension.
@@ 3.1 Sparse Learning for Ising Model Reconstruction
We begin by introducing notations for the presentation of our framework. Let ${J}_{i}$ be the $i$-th column of the interaction matrix ${J}$, then for each node $i$ in this network, we define logarithmic pseudo-likelihood (PL) as
${\mathcal{L}}_{n}({J}_{i})\coloneqq$ $\sum_{r=1}^{n}\log P({}_{i}^{(r)}|\ {}_{1}^{(r)},\ldots,{}_{i-1}^{(r)},{}_{i+1}^{(r)},\ldots,{}_{p}^{(r)}).$
@@
The advantage of SLIDE is two-fold. First, it meets the need for reconstructing networks by selecting the neighbors of each node. Once neighbors of every node are correctly found, the network structure must be identical to the true one. Particularly, the local system complexity constraint conducts the selection of neighbors straightforwardly, which is the key reason for the superiority of SLIDE. In contrast, the other indirect constraints would need higher sample complexity to implicitly select the neighbors. Second, it exploits the PL $\widehat{{J}}_{i}\;(i=1,\ldots,p)$ as a surrogate of the computationally intractable likelihood function. It has an explicit expression:
${\mathcal{L}}_{n}({J}_{i})=-\sum_{r=1}^{n}\log\Big{\{}\exp\Big{(}-2\sum_{j:j\neq i}{J}_{ij}{}_{i}^{(r)}{}_{j}^{(r)}\Big{)}+1\Big{\}},$
and computing PL has an $O(np)$ time complexity and thus can be quickly computed.
@@
A careful reader may notice that (4) does not exploit the symmetry of ${J}$ to learn the structure. On the one hand, this is because the solution of (4) already exactly recover the underlying graph structure (see Theorem 1). On the other hand, in the asymptotic regime, accounting for the symmetry of ${J}$ does not offer additional benefits to the recovery but causes the computational process to be more complicated as it has to consider $O(p^{2})$ parameters simultaneously. In the finite-sample regime, we acknowledge leveraging symmetry might improve numerical performance (Höfling and Tibshirani, 2009; Miasojedow and Rejchel, 2018), and we develop a symmetry-aware SLIDE variant in Supplementary Materials.
@@ 3.2 Sample Complexity of SLIDE
Allowing the quadruple $({\lambda},{\gamma},d,p)$ to scale with the sample size $n$, the following theorem gives the theoretic sample complexity of SLIDE to successfully recover the underlying graph structure.
@@
Assume that $e^{-2{\gamma}}{\lambda}^{2}\leq 8$. Suppose that $z^{(1)},\ldots,z^{(n)}$ are i.i.d drawn from some $P_{{J}^{*}}\in\mathcal{G}_{p,d}({\lambda},{\gamma})$. For any constant $\varepsilon\in(0,1)$, if the sample size $n$ satisfies
@@
Our sample complexity $n^{*}$ can be presented more concisely as
@@
Clearly, the sample complexity of SLIDE lies in the globally optimal regime (3), showing SLIDE is a globally optimal estimator. Furthermore, to the best of our knowledge, along every axis of $({\lambda},{\gamma},d,p)$, it matches or improves the theoretic sample complexities of all existing state-of-the-art algorithms listed in Table 1.
@@
Projection-type estimator (Santhanam and Wainwright, 2012) $e^{4{\gamma}}$ $d^{2}$ $\log p$ ${\lambda}^{-2}$
$\ell_{1}$-constrained nodewise regression (Rigollet and Hütter, 2017) $e^{8{\gamma}}$ $d^{2}$ $\log p$ ${\lambda}^{-2}$
@@
$\ell_{1}$-regularized approximated MLE (Miasojedow and Rejchel, 2018) $\geq e^{4{\gamma}}$ $d^{2}$ $\log p$ ${\lambda}^{-2}$
@@
    Miasojedow and Rejchel (2018) does not explicitly demonstrate how sample complexity depends on ${\gamma}$. In our analysis, we derive a dependency of $\geq e^{4{\gamma}}$ based on their reported sample complexity under a specific coupling matrix configuration. For further details, please refer to Supplementary Materials.
@@
    The comparison to a similar estimator proposed by Raskutti et al. (2011) is discussed in Supplementary Materials.
The SLIDE’s sample complexity scales linearly in the maximum number of neighbors $d$, while other existing methods have theoretic sample complexities growing quadratically in $d$. This is particularly important in practice, as systems in fields such as neuroscience and sociology often involve networks with a larger $d$ (Grabowski, 2009; Marinazzo et al., 2014). In these domains, our method would demonstrate superior performance compared to existing algorithms. Theoretically, our method allows the true sparse network to have a higher node degree. Most notably, our method proves a conjecture in Santhanam and Wainwright (2012) as its sample complexity scales linearly with $d$. We know that the optimal sample complexity scales with $d^{c}$, where $c\in[1,2]$ is an unknown precise number conjectured to be $1$, according to the work of Santhanam and Wainwright (2012). By providing support for the claim that $r=1$ for the SLIDE, our finding verifies the truth of this supposition.
For the maximum of the maximum neighborhood weight ${\gamma}$, the SLIDE also lies in the globally optimal regime. In the case of large ${\gamma}$, the underlying Ising model is low-temperature, which is shown to be challenging to reconstruct (Montanari and Pereira, 2009; Nguyen and Berg, 2012) and attracts keen attention from researchers (Decelle and Ricci-Tersenghi, 2014; Nguyen and Berg, 2012). Roughly speaking, the difficulty arises since in this case, most samples we collect would be of the network’s ground states. The sample size needed to include a sufficient number of non-trivial samples grows exponentially as ${\gamma}$ increases. Therefore, SLIDE has a greater potential for widespread use because it uses significantly fewer samples to reconstruct a low-temperature Ising model than existing approaches. Examples illustrating this advantage are provided in Figure 2.
Last but not least, our proposed sample complexity for a given number of nodes $p$ and the minimum interaction ${\lambda}$ is a perfect match to the optimal sample complexity given by Santhanam and Wainwright (2012), i.e., it is proportional to $\log{p}$ and ${\lambda}^{-2}$. For high-dimensional binary variables, where $p$ can scale exponentially with $n$, the former result is significant. The last implication means that no alternative structure recovery approach can use fewer samples than our proposal when the signal-to-noise ratio is relatively low.

### v3

@@ Abstract
The reconstruction of interaction networks between random events is a critical problem arising from statistical physics and politics, sociology, biology, psychology, and beyond. The Ising model lays the foundation for this reconstruction process, but finding the underlying Ising model from the least amount of observed samples in a computationally efficient manner has been historically challenging for half a century. Using sparsity learning, we present an approach named SLIDE whose sample complexity is globally optimal. Furthermore, an algorithm is developed to give a statistically consistent solution of SLIDE in polynomial time with high probability. On extensive benchmarked cases, the SLIDE approach demonstrates dominant performance in reconstructing underlying Ising models, confirming its superior statistical properties. The application on the U.S. senators voting in the six congresses reveals that both the Republicans and Democrats noticeably assemble in each congress; interestingly, the assembling of Democrats is particularly pronounced in the latest congress.
Keywords: Ising model, sparse learning, sample complexity, polynomial time complexity, consistent algorithm
@@ 1 Introduction
In contemporary applications, the Ising model serves as a prevalent tool for characterizing the pairwise interactions between binary variables in a network system. In this network, each node represents a binary variable, the presence of an edge between any two nodes indicates that the corresponding variables have an interaction. Reconstructing the corresponding graph of the Ising model is critical to the studies of interaction in a broad spectrum of complex systems in biology, neuroscience, ecology, economics, political science, image analysis (Cocco et al., 2009; Weistuch et al., 2021; Noble et al., 2018; Battiston et al., 2016; Banerjee et al., 2008; Saremi & Sejnowski, 2013). While many methods have been proposed during the past half century for reconstructing Ising models, finding the most sample-efficient estimate in a computationally efficient manner such as in a polynomial time remains an open problem. Given the profound importance of the Ising model, especially in the era of social media, there is an urgent need to solve this open problem.
@@ 1.1 Literature Review
Reconstructing Ising models by maximizing the likelihood function is considered to be the gold standard, although it is computationally intractable since the evaluation of the partition function involves the sum of $2^{p}$ terms (Nguyen et al., 2017), where $p$ is the number of nodes in the Ising model. Many techniques have been developed to approximate the likelihood function. For instance, it is suggested that Monte Carlo methods can be used to approximate the partition function (Geyer, 1994; Miasojedow & Rejchel, 2018), and then the approximating likelihood function is maximized. Although this approach is asymptotically exact, it takes an exponential number of runs to achieve a predefined accuracy. Alternative methods that can avoid evaluating the partition function include the naive mean-field approximation and its advanced variants (Chow & Liu, 1968; Kappen & de Borja Rodríguez Ortiz, 1997; Nguyen & Berg, 2012). These methods assume that the network is composed of some basic clusters, and when this assumption is met, they give the exact value of the likelihood function. As one increases the size of basic clusters, the methods apply to a broader class of networks, while the computational complexity also sharply grows. However, for networks beyond the applicable class, the effectiveness of such approximations remains uninvestigated.
@@
Instead of approximations, surrogates of the likelihood function also have been proposed to sidestep the computational cost. Ravikumar et al. (2010) proposed to node-wisely maximize the conditional likelihood. The product of conditionals, commonly referred to as pseudo-likelihood or composite likelihood, is often coupled with various regularizations and widely applied in the literature (Besag, 1975; Höfling & Tibshirani, 2009; Jalali et al., 2011; Xue et al., 2012). In Vuffray et al. (2016), one proposed to use the interaction screening objective, which is closely related to the exponential loss of the logistic model. These methods are capable for general Ising models and can alleviate the prohibitive computation for their reconstruction. Typically, this strategy employs convex regularization, which may introduce bias and the subtle steps in tuning hyperparameters and selecting post-inference thresholds for theoretic guarantees (Lokhov et al., 2018). More importantly, we prefer methods that use the fewest samples to perfectly recover the underlying Ising model. However, every aforementioned method requires a sample size much larger than the necessary sample complexity derived in Santhanam & Wainwright (2012). The best result is the optimality along only a subset of the parameters (Lokhov et al., 2018).
@@
Graphical structure recovery resides in the realm of variable selection, for which the predominance of sparse learning approach has been demonstrated in a variety of statistical models (Raskutti et al., 2011; Sara van de Geer & Peter Bühlmann, 2013). The sparse learning approach directly addresses the biasedness of the aforementioned regularization-based methods. Moreover, the approach may lead to stronger statistical properties, as they usually result in a smaller parameter space to search over. Finally, the development of statistical tools allows us to consider directly solving the sparse learning problem. For some well-studied models like (generalized) linear models, in recent years, statisticians have proposed several feasible approximate algorithms that solve the sparse regression under mild conditions (Huang et al., 2018; Zhu et al., 2020; Zhang et al., 2023; Zhu et al., 2022). However, for Ising models, the potential benefits of the sparse learning approach have not yet been rigorously authenticated, and how to design a feasible algorithm that captures the essence of sparse learning for reconstructing the Ising model remains another meaningful problem.
@@ 2.1 Ising Model and its Reconstruction
Definition 1 (Family of distributions, Santhanam & Wainwright (2012))
@@
[…]
since ${J}_{ij}=0$ implies $(i,j)\notin E$. Note that, we occasionally omit the subscript J{J} in $P_{{J}}$ for simplicity when there is no ambiguity; besides, we define the support set of matrix J{J} as ${\mathrm{supp}}({J})$.
Suppose ${J}^{*}$ is the unknown true parameters that characterizes variables’ interactions, Ising model reconstruction aims to recover ${J}^{*}$ via $n$ observed samples $z^{(1)},\ldots,z^{(n)}\in\mathbb{R}^{p}$ that are independently sampled from the Ising model $P_{{J}^{*}}$. Specifically, in this paper, we focus on this problem when the underlying graph only has a relatively small number of edges, i.e., the vast majority of the ${J}_{ij}$’s are exactly zero. This sparsity assumption is reasonable as it gives better interpretability of the model. In the meanwhile, it appears to be a necessary assumption when the dimension $p$ is comparable to or larger than the sample size $n$ (the so-called high-dimensional regime).
@@ 2.2 Globally Optimal Sample Complexity Regime
Background. For the reconstruction of Ising models, the sample complexity of an estimator refers to the minimum sample size such that, with a high probability, this estimator can recover any underlying Ising model belongs to the family of distributions introduced in Definition 1. For Ising models satisfying Definition 1, Santhanam & Wainwright (2012) derived the information-theoretic lower bound of the sample size with respect to four structural parameters: maximum neighborhood weight $\gamma$}, maximum node degree $d$, number of nodes $p$, and minimum signal $\lambda$. These parameters are of primary interest when studying the sample complexity. Specifically, Santhanam & Wainwright (2012) showed that the information-theoretic lower bound is of order
@@
Moreover, to show the tightness of the above lower bound, Santhanam & Wainwright (2012) provided an attainable sample complexity, by analyzing the sample complexity of a projection-type graphical structure estimator. To be specific, this estimator returns the graphical structure of the model whose second-order moments best match their empirical counterparts. Such an estimator is computationally intractable, but it offers a sample complexity that is attainable (at least theoretically) by an estimator:
@@
The optimal sample size $n_{\text{opt}}$ lies within $n_{l}$ and $n_{u}$ yet its specific form remains unknown. From $n_{l}$ and $n_{u}$, we know that the optimal sample complexity $n_{\text{opt}}$ has the following dependencies on $({\lambda},{\gamma},d,p)$: $n_{\text{opt}}\propto\log p$, $n_{\text{opt}}\propto{\lambda}^{-2}$, $n_{\text{opt}}\propto d^{r}$ and $n_{\text{opt}}\propto e^{c{\gamma}}$ where $r$ and $c$ are two constants satisfying $r\in[1,2]$ and $c\in[1,4]$.
@@
for any $c\in[1,4]$. An estimator whose sample complexity satisfies (3) is called a globally optimal estimator. The regime (3) integrates insights from both the information-theoretic lower bounds (1) and the attainable sample complexities (2) derived by Santhanam & Wainwright (2012). By defining the regime as the lower and upper limits for each critical structural parameter, this regime comprehensively covers the entire range of these parameters. The rationality of the globally optimal regime (3) is twofold.
  • Tightness. To the best of our knowledge, no existing algorithm has a lower sample complexity than our upper boundary, showing this regime as the tightest one that contains the optimal rate in the literature. Additionally, this regime incorporates the conjecture by Santhanam & Wainwright (2012) that the information-theoretic lower bound for $d$ is tight, further reinforcing the tightness of the defined regime.
@@
    overlooked the dependency on the maximum degree $d$. Their regime implicitly assumes $d=o(e^{\gamma})$, where the effect of $d$ in both of (1) and (2) can be ignored, and they merely focus on the effect of $\gamma$}. However, this assumption may not hold for general Ising models and a counterexample is constructed in Section 5.1.1. By removing the implicit assumption, the globally optimal regime serves as a more comprehensive extension.
@@ 3.1 Sparse Learning for Ising Model Reconstruction
We begin by introducing notations for the presentation of our framework. Let ${J}_{i}$ be the $i$-th column of the interaction matrix J{J}, then for each node $i$ in this network, we define logarithmic pseudo-likelihood (PL) as
${\mathcal{L}}_{n}({J}_{i})\coloneqq$ $\sum_{r=1}^{n}\log P(\real_{i}^{(r)}|\ \real_{1}^{(r)},\ldots,\real_{i-1}^{(r)},\real_{i+1}^{(r)},\ldots,\real_{p}^{(r)}).$
@@
The advantage of SLIDE is two-fold. First, it meets the need for reconstructing networks by selecting the neighbors of each node. Once neighbors of every node are correctly found, the network structure must be identical to the true one. Particularly, the local system complexity constraint conducts the selection of neighbors straightforwardly, which is the key reason for the superiority of SLIDE. In contrast, the other indirect constraints (Lokhov et al., 2018) would need higher sample complexity to implicitly select the neighbors. Second, it exploits the PL ${\mathcal{L}}_{n}({J}_{i})\;(i=1,\ldots,p)$ as a surrogate of the computationally intractable likelihood function. It has an explicit expression:
${\mathcal{L}}_{n}({J}_{i})=-\sum_{r=1}^{n}\log\Big\{\exp\Big(-2\sum_{j:j\neq i}{J}_{ij}\real{}_{i}^{(r)}\real{}_{j}^{(r)}\Big)+1\Big\},$
and computing the PL has an $O(np)$ time complexity and thus can be efficiently computed.
@@
A careful reader may notice that (4) does not exploit the symmetry of J{J} to learn the structure. On the one hand, this is because the solution of (4) already exactly recover the underlying graph structure (see Theorem 1). On the other hand, in the asymptotic regime, accounting for the symmetry of J{J} does not offer additional benefits to the recovery but causes the computational process to be more complicated as it has to consider $O(p^{2})$ parameters simultaneously. In the finite-sample regime, we acknowledge leveraging symmetry might improve numerical performance (Höfling & Tibshirani, 2009; Miasojedow & Rejchel, 2018), and we develop a symmetry-aware SLIDE variant in the Supplementary Materials.
@@ 3.2 Sample Complexity of SLIDE
Allowing the quadruple $({\lambda},{\gamma},d,p)$ to scale with the sample size $n$, the following theorem gives the theoretic sample complexity of SLIDE to successfully recover the underlying graphical structure.
@@
Assume that $e^{-2{\gamma}}{\lambda}^{2}\leq 8$. Suppose that $z^{(1)},\ldots,z^{(n)}$ are i.i.d. drawn from some $P_{{J}^{*}}\in\mathcal{G}_{p,d}({\lambda},{\gamma})$. For any constant $\varepsilon\in(0,1)$, if the sample size $n$ satisfies
@@
Our sample complexity $n^{*}$ can be presented more concisely (up to some logarithmic factors) as
@@
Clearly, the sample complexity of SLIDE lies in the globally optimal regime (3), showing that SLIDE is a globally optimal estimator. Furthermore, to the best of our knowledge, along every axis of $({\lambda},{\gamma},d,p)$, it matches or improves the theoretic sample complexities of all existing state-of-the-art algorithms listed in Table 1.
@@
Projection-type estimator (Santhanam & Wainwright, 2012) $e^{4{\gamma}}$ $d^{2}$ $\log p$ ${\lambda}^{-2}$
$\ell_{1}$-constrained nodewise regression (Rigollet & Hütter, 2017) $e^{8{\gamma}}$ $d^{2}$ $\log p$ ${\lambda}^{-2}$
@@
$\ell_{1}$-regularized approximated MLE (Miasojedow & Rejchel, 2018) $\geq e^{4{\gamma}}$ $d^{2}$ $\log p$ ${\lambda}^{-2}$
@@
    Miasojedow & Rejchel (2018) does not explicitly demonstrate how sample complexity depends on $\gamma$}. In our analysis, we derive a dependency of $\geq e^{4{\gamma}}$ based on their reported sample complexity under a specific coupling matrix configuration. For further details, please refer to the Supplementary Materials.
@@
    The comparison to a similar estimator proposed by Raskutti et al. (2011) is discussed in the Supplementary Materials.
The SLIDE’s sample complexity scales linearly in the maximum number of neighbors $d$, while other existing methods have theoretic sample complexities growing quadratically in $d$. This is particularly important in practice, as systems in fields such as neuroscience and sociology often involve networks with a larger $d$ (Grabowski, 2009; Marinazzo et al., 2014). In these domains, our method would possess superior performance compared to existing algorithms. Theoretically, our method allows the true sparse network to have a higher node degree. Most notably, our method proves a conjecture in Santhanam & Wainwright (2012) as its sample complexity scales linearly with $d$. We know that the optimal sample complexity scales with $d^{r}$, where $r\in[1,2]$ is an unknown precise number conjectured to be $1$, according to the work of Santhanam & Wainwright (2012). By providing support for the claim that $r=1$ for the SLIDE, our finding verifies the truth of this supposition.
For the maximum neighborhood weight $\gamma$}, the SLIDE also lies in the globally optimal regime. In the case of large $\gamma$}, the underlying Ising model is low-temperature, which is shown to be challenging to reconstruct (Montanari & Pereira, 2009; Nguyen & Berg, 2012) and attracts keen attention from researchers (Decelle & Ricci-Tersenghi, 2014; Nguyen & Berg, 2012). Roughly speaking, the difficulty arises since in this case, most samples we collect would be of the network’s ground states. To ensure a sufficient number of non-trivial samples, the sample size needs to grow exponentially as $\gamma$} increases. Therefore, SLIDE has a greater potential for widespread use because it requires significantly fewer samples to reconstruct a low-temperature Ising model than existing approaches. Examples illustrating this advantage are provided in Figure 2.
Last but not least, the sample complexity of SLIDE in terms of the number of nodes $p$ and the minimum interaction $\lambda$} is a perfect match to the optimal sample complexity given by Santhanam & Wainwright (2012), i.e., it is proportional to $\log{p}$ and ${\lambda}^{-2}$. For high-dimensional binary variables, where $p$ can scale exponentially with $n$, the former result is significant. The last implication means that no alternative structure recovery approach can use fewer samples than our proposal when the signal-to-noise ratio is relatively low.